OpenStudy (anonymous):

Consider the function f(x)=3x^2 – 5x on the interval [-4,4]. Find the average slope of the function on this interval. By the mean value theorem, we know there exists at least one c in the open interval (-4,4) such that f’(c) is equal to this mean slope. For this problem, there are 2 values of c that work. The smaller one and the larger one are?

OpenStudy (amistre64):

f(4)- f(-4) --------- = average slope of the interval 4--4

OpenStudy (amistre64):

determine when f'(x) = that value and youll have your cs

OpenStudy (anonymous):

i got slope of -5 and c is 0 but that isn't right..

OpenStudy (amistre64):

f' = 6x-5 -5 = 6x-5 when x=0 why isnt that right?

OpenStudy (amistre64):

this might be one of those mean integral things with 1/(b-a) then

OpenStudy (amistre64):

we sure there aint a "^3" in this?